OpenStudy (anonymous):

A bicyclist is moving at a speed of 5.5 m/s. she squeezes the brakes, giving her an acceleration of -2.1 m/s^2 . how fast the biciclist after 0.3 seconds of applying the brakes.

OpenStudy (isaidavila):

Vf=Vo-at Vf=5.5-(2.1)(0.3) Vf=??